OpenStudy (anonymous):

4x^-1/(4x)^-2 need to simplify the expression

OpenStudy (angela210793):

=64x

OpenStudy (anonymous):

how.

OpenStudy (angela210793):

(4/x) ---------=4/x*16x^2=4*168x=64x (1/16x^2)

OpenStudy (angela210793):

4*16*x=64x sorry shift wasn't working :/

OpenStudy (anonymous):

so. confused. lol

OpenStudy (angela210793):

\[x ^{-n}=1/x ^{n}\] tht's wht i used

OpenStudy (anonymous):

well yeyah so far I have 4/x/1/16x^2

OpenStudy (anonymous):

I think there is a trick here I don't know how to do.

OpenStudy (angela210793):

(a/b)/(c/d)=a/b*d/c right? this is wht i' used